OpenStudy (anonymous):

xdy+ydx=(x^3y^6)dx answer: 1/y^5=-5x^3/2+(c1(x^5))

OpenStudy (anonymous):

1.divide by dx 2.divide by xy^6 3. put \[y ^{-5}=t,-5y ^{-6}\frac{ dy }{ dx }=\frac{ dt }{ dx }\]

OpenStudy (anonymous):

Another approach (though essentially the same as @Princer_Jones's ): \[\begin{align*} xy'+y&=x^3y^6&\text{set }t=yx\\\\ t'&=t^3y^3\\\\ t'&=\frac{t^6}{x^3} \end{align*}\]

OpenStudy (anonymous):

i am trying to understand

OpenStudy (anonymous):

@rubcksy there will not be any negative in the rhs..

OpenStudy (anonymous):

\[y ^{-6}\frac{ dy }{ dx }+\frac{ 1 }{ x }y ^{-5}=x^2\] put \[y ^{-5}=t,-5y ^{-6}\frac{ dy }{ dx }=\frac{ dt }{ dx }\] \[y ^{-6}\frac{ dy }{ dx }=-\frac{ 1 }{ 5 }\frac{ dt }{ dx }\] \[-\frac{ 1 }{ 5 }\frac{ dt }{ dx }+\frac{ 1 }{ x }t=x^2\] \[\frac{ dt }{ dx }+\frac{ -5 }{ x }t=-5x^2\] \[I.F=e ^{\int\limits \frac{ -5 }{ x }}dx=e ^{-5\ln x}=e ^{\ln x ^{-5}}=x ^{-5}\] C.S. is \[tx ^{-5}=\int\limits x^2x ^{-5}dx=\int\limits x ^{-3}dx+c\] complete it.